A man buys 1000 government bonds. Given that the yield on bond is 20% per annum , current and expected interest rates are 5% and 2.5% respectively. Determine the current and expected price of bond, capital gains, total expected yield on bond , and critical rate of interest
1
Expert's answer
2021-08-16T17:33:12-0400

No. of bonds =1000

Yield on bond = 20%

Current interest rate =5%

Expected interest rate=2.5%

Current price of bond = "1000\times \frac {5}{100}+1000"

= 1050.

Expected price of bond = "1000\times \frac {2.5}{100}+1000"

=1025.

Total expected yield on bond "= 1050-1025"

"=25"